OpenStudy (anonymous):

solve by elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (retireed):

I don't no if this is the best way to solve this, but it worked. A better way to do this would involve creating a matrix and solving the determinants, but I haven't done that since college. I solved this by subtracting the second equation from the first one. This wii eliminate the x term and leave an equation in terms of y and z, call it equation 4. Then I added equation 1 and 3 together, again eliminating x, creating equation 5. I then multiplied equation 4 by the number 3 and subtracted the new equation from equation 5. The result will eliminate the z term and solve for y. Then plug y into into equation 4 or 5 and solve for z. Then plug y and z into equation 1 2 or 3 and solve for x. The answers for x, y and z might seem wrong, but if you double check them into all three equations you can verify their validity.

OpenStudy (dayakar):

can u show your work

OpenStudy (dayakar):

-2x+2y+3z =0--------(1) -2x-y+z = -3---------(2) 2x+3y+3z=5-------(3)

OpenStudy (dayakar):

add (1) +(3) u can get 5y+6z = 5-------(4) (2) +(3) 2y +4z =2-------------(5)

OpenStudy (dayakar):

2* (4) ------ 10y+12z=10 ------(6) 5*(5)------ 10y+20z =10 -------(7) subtract (6) from (7) 8z =0 z= 0-----(8) plug z=0 in (6) you can get y= 1 plug z=0 and y=1 in (1) you can get x=1 therefore x=1, y=1 and z=0

OpenStudy (denisaboichuk):

z=0